OpenStudy (anonymous):

cos x - 5x = 4 sin 3x sin x cos x

OpenStudy (dumbcow):

http://en.wikipedia.org/wiki/Trig_identities#Power-reduction_formula \[\cos x - \cos(5x) = 2 \sin(3x) \sin(2x)\]